OpenStudy (baby456):

5. Solving equations by completing the square! r^2+2r-33=0. 6. a^2-2a-48=0

OpenStudy (baby456):

@wolf1728

OpenStudy (baby456):

please help me!

OpenStudy (anonymous):

Solve Quadratic Equation by Completing The Square 2.2 Solving r2+2r-33 = 0 by Completing The Square . Add 33 to both side of the equation : r2+2r = 33 Now the clever bit: Take the coefficient of r , which is 2 , divide by two, giving 1 , and finally square it giving 1 Add 1 to both sides of the equation : On the right hand side we have : 33 + 1 or, (33/1)+(1/1) The common denominator of the two fractions is 1 Adding (33/1)+(1/1) gives 34/1 So adding to both sides we finally get : r2+2r+1 = 34 Adding 1 has completed the left hand side into a perfect square : r2+2r+1 = (r+1) • (r+1) = (r+1)2 Things which are equal to the same thing are also equal to one another. Since r2+2r+1 = 34 and r2+2r+1 = (r+1)2 then, according to the law of transitivity, (r+1)2 = 34 We'll refer to this Equation as Eq. #2.2.1 The Square Root Principle says that When two things are equal, their square roots are equal. Note that the square root of (r+1)2 is (r+1)2/2 = (r+1)1 = r+1 Now, applying the Square Root Principle to Eq. #2.2.1 we get: r+1 = √ 34 Subtract 1 from both sides to obtain: r = -1 + √ 34 Since a square root has two values, one positive and the other negative r2 + 2r - 33 = 0 has two solutions: r = -1 + √ 34 or r = -1 - √ 34

OpenStudy (wolf1728):

r^2+2r-33=0 To complete the square: Move the "non X" term to the right: r^2+2r=33 Divide the equation by the coefficient of X² which in this case is 1 so there's no change r^2+2r=33 Now for "completing the square" we: take the coefficient of X which is 2 divide that by 2 1 square that number 1 then add it to both sides of the equation r^2+2r +1=34 Take the square root of BOTH sides of the equation (r+1) = sqrt(34) r = sqrt(34) -1 site explaining completing the square: http://www.1728.org/quadr2.htm

OpenStudy (anonymous):

Solving a2-2a-48 = 0 by Completing The Square . Add 48 to both side of the equation : a2-2a = 48 Now the clever bit: Take the coefficient of a , which is 2 , divide by two, giving 1 , and finally square it giving 1 Add 1 to both sides of the equation : On the right hand side we have : 48 + 1 or, (48/1)+(1/1) The common denominator of the two fractions is 1 Adding (48/1)+(1/1) gives 49/1 So adding to both sides we finally get : a2-2a+1 = 49 Adding 1 has completed the left hand side into a perfect square : a2-2a+1 = (a-1) • (a-1) = (a-1)2 Things which are equal to the same thing are also equal to one another. Since a2-2a+1 = 49 and a2-2a+1 = (a-1)2 then, according to the law of transitivity, (a-1)2 = 49 We'll refer to this Equation as Eq. #3.2.1 The Square Root Principle says that When two things are equal, their square roots are equal. Note that the square root of (a-1)2 is (a-1)2/2 = (a-1)1 = a-1 Now, applying the Square Root Principle to Eq. #3.2.1 we get: a-1 = √ 49 Add 1 to both sides to obtain: a = 1 + √ 49 Since a square root has two values, one positive and the other negative a2 - 2a - 48 = 0 has two solutions: a = 1 + √ 49 or a = 1 - √ 49

OpenStudy (baby456):

i really don't get it!

OpenStudy (baby456):

i am so confused all i see are numbers!

OpenStudy (baby456):

@ganeshie8

OpenStudy (wolf1728):

Gee I got bounced out of here a while ago and I see cutepochacco finished the OTHER question!!! Wow - that is fast !! So you want some explanation baby456?

OpenStudy (baby456):

yes please

OpenStudy (wolf1728):

Gee I can't really explain why completing the square works

OpenStudy (baby456):

its ok ! i guess i will watch some youtube videos!

OpenStudy (wolf1728):

okay - they might tell you the how but not WHY it works.

OpenStudy (baby456):

its okay i will learn! bye!

OpenStudy (wolf1728):

okay good luck :-)